I need help with these Laplace problems:)

(1 point) Find the Laplace transform of <9 f(t) = { 0, I(t - 9)?, 129 F(s) =

(1 point) Find the inverse Laplace transform of e-75 F(s) = 52 – 2s – 15 f(t) = . (Use step(t-c) for uc(t).)

(1 point) Find the Laplace transform of 0. f(t) t<5 112 – 10t + 30, 125 F(s) =
Answer #1

txa & flt) = so t = {(+-9² +39 Plt) = (4-9)”. ult-9) 1{f(}- 43(4-93 ult-933 1715)=21 e-75 * F182 - Les MS F (S) 25-15 L{ ec#16) = o stas 1 E-106+30 +3,5 flt) (-10t+30) ult-5) flt) t-5)45] ult-5) f(t). E-5)+-5) + sult-5) Lf4lt)}-{{6-5)*ult-s)}+2{sul